OpenStudy (anonymous):

Antiderivative tan^2x/x^2+1

OpenStudy (trojanpoem):

\[\int\limits_{}^{}\frac{ \tan^2x }{ x^2 + 1 } dx\]

OpenStudy (trojanpoem):

This ?

OpenStudy (trojanpoem):

or \[\int\limits_{}^{}\frac{ (\tan^{-1}x)^2 }{ x^2 + 1 } dx\]

OpenStudy (trojanpoem):

for the second : \[\int\limits_{}^{} (\tan^{-1}x)^2 dtan^{-1}x = \frac{ 1 }{ 3 }(\tan^{-1}x)^3 + c\]

OpenStudy (trojanpoem):

for the first: IDK.